THE ULTRA VIOLET SPECTUM OF THE NCN RADICAL

\underline{S. A. BEATON}, J. M. BROWN , Physical and Theoretical Chemistry Laboratory, South Parks Road, Oxford, UK, OX1 3QZ ; Y. ITO , Department of Mechanical Engineering, Nagaoka University of Technology, 1603-1 Kamitomioka, Nagaoka, Niigata 940-21, Japan .

The (0,0) band of the Ã^{3}Π_{u} − X̃^{3}Σ_{g}^{−} transition of NCN, observed previously by Herzberg and Travis^{a}, has been reinvestigated by means of laser-induced fluorescence at 329 nm. A frequency doubled dye laser was used to record the spectrum of NCN downstream of a microwave discharge through N_{2} and CF_{4} in He^{b}. From the rotational analysis, improved values for the molecular constants of NCN in both states were obtained, including Λ-type doubling parameters for the ^{3}Π_{u} upper state. The main molecular constants obtained were

X̃^{3}Σ_{g}^{−}: B′′_{000} = 0.3972458(95)cm^{−1} λ′′ = 0.79402(50)cm^{−1}
Ã^{3}Π_{u}: B′_{000} = 0.3966487(95)cm^{−1} A′ =–37.4493(38)cm^{−1}
T′_{000} = 30383.8110(51)cm^{−1}

The B values (and hence bond lengths) are very similar in the two states and therefore the spectrum is dominated by Δv = 0 transitions. The (010)-(010) band occurs in a similar wavelength region, the excited state of this transition exhibiting the Renner-Teller effect. Analyses of both the ^{3}Δ_{g} −^{3} Π_{u} and the ^{3}Σ_{g} −^{3} Π_{u} components of this hot band will be presented.

aG.HERZBERG AND D.N.TRAVIS, Can.J.Phys.42,1658-1675 (1964).

bG.P.SMITH, R.A.COPELAND AND D.R.CROSLEY, J.Chem.Phys.91,1987-1993 (1989).